# Destructive and constructive RIS beamforming in an ISAC-multi-user MIMO network

### v1

@@ III-B RIS phase-shift optimization
We now move on to the second part of the AO where we obtain $\boldsymbol{\Theta}$ while $\mathbf{F}$ is fixed. We rewrite $\rho_{l}$, defined in (6), as
@@
$\quad+2\Re\left(\text{Tr}\left(\delta_{m}^{*}\boldsymbol{\Theta}\mathbf{T}^{\mathsf{H}}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{H}_{s,{\rm T}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\right)\right)$
$\quad+\delta_{s}^{2}\text{Tr}\left(\mathbf{H}_{s,{\rm T}}^{\mathsf{H}}\mathbf{H}_{s,{\rm T}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\right).$ (12)
Utilizing the identity $\text{Tr}(\mathbf{A}\mathbf{X}\mathbf{B}\mathbf{X})=\left(\text{vec}\left(\mathbf{X}^{\top}\right)^{\top}\right)\left(\mathbf{B}^{\top}\otimes\mathbf{A}\right)\text{vec}(\mathbf{X})$ and $\mathbf{T}=\mathbf{T}^{\mathsf{H}}$, the first term of $\rho_{l}$ can be rewritten as
$\text{vec}\left(\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}}\right)^{\mathsf{H}}\left(\left(\mathbf{H}_{t}^{*}\mathbf{H}_{t}^{\top}\right)\otimes\left(\delta_{r}^{2}\mathbf{H}_{\text{t}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\right)\right)\text{vec}\left(\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}}\right).$ (13)
Then, thanks to the nature of $\boldsymbol{\Theta}$ [10] we can write that $\text{vec}(\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}})^{\mathsf{H}}=\text{diag}(\text{vec}(\mathbf{T}))^{\mathsf{H}}\tilde{\boldsymbol{\theta}}$, where $\tilde{\boldsymbol{\theta}}$ collects the diagonal entries of $\boldsymbol{\Theta}^{*}\otimes\boldsymbol{\Theta}$. We can finally rewrite the first term of $\rho_{l}$ as $\tilde{\boldsymbol{\theta}}^{\mathsf{H}}\mathbf{Q}_{l}\tilde{\boldsymbol{\theta}}$, where
$\mathbf{Q}_{l}$ $=\left(\text{diag}\left(\text{vec}\left(\mathbf{T}\right)\right)\right)^{\mathsf{H}}\left(\left(\mathbf{H}_{t}^{*}\mathbf{H}_{t}^{\top}\right)\otimes\left(\delta_{r}^{2}\mathbf{H}_{\text{t}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\right)\right)$
@@
$\mathbf{p}_{l}$ $=\text{diag}(\text{vec}(\mathbf{T}))^{\mathsf{H}}\text{vec}\left(\delta_{m}^{*}\mathbf{H}_{t}\mathbf{H}_{s,{\rm T}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\right).$ (16)
@@
$=\frac{1}{\sigma_{\rm T}^{2}}\sum_{l=1}^{L}2\Re\left(\boldsymbol{\theta}^{\mathsf{H}}\left(\Sigma\left(\mathbf{b}_{l}^{(t)}\right)+\Sigma\left(\mathbf{p}_{l}\right)\right)\boldsymbol{\theta}\right),$ (18)
where $\Sigma\left(\mathbf{b}_{l}^{(t)}\right)\in\mathbb{C}^{N\times N}$ is a matrix built by stacking the elements of $\mathbf{b}_{l}^{(t)}$. We define the variable $\hat{\boldsymbol{\Theta}}=\hat{\boldsymbol{\theta}}\hat{\boldsymbol{\theta}}^{\mathsf{H}}$, where $\hat{\boldsymbol{\theta}}=[\boldsymbol{\theta}^{\top},1]^{\top}$ and $\mathbf{H}_{r,k}=\text{diag}(\mathbf{h}_{r,k})$. Then, the second subproblem of the AO algorithm is
@@
$\text{Tr}\left(\mathbf{R}_{k,k}\hat{\boldsymbol{\Theta}}\right)\geq\gamma_{k}\left(\sum_{l\neq k}\text{Tr}\left(\mathbf{R}_{k,l}\hat{\boldsymbol{\Theta}}\right)+\sigma^{2}_{k}\right),$ (19c)
@@
where
$\mathbf{R}_{k,l}$ $=\begin{bmatrix}\mathbf{H}_{r,k}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\mathbf{H}_{r,k}&\mathbf{H}_{r,k}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{h}_{s,k}\\ \mathbf{h}_{s,k}^{\mathsf{H}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{\text{t}}^{\mathsf{H}}\mathbf{H}_{r,k}&|\mathbf{h}_{s,k}^{\mathsf{H}}\mathbf{f}_{l}|^{2}\end{bmatrix},$ (20)
$\boldsymbol{\Xi}_{l}^{(t)}$ $=\begin{bmatrix}\Sigma\left(\mathbf{b}_{l}^{(t)}\right)+\Sigma(\mathbf{p}_{l})&\mathbf{0}_{N\times 1}\\ \mathbf{0}_{1\times N}&0\end{bmatrix}.$ (21)
Once again by removing the rank-one constraints, the relaxed problem becomes convex. Hence, it can be solved by CVX and we can retrieve a good sub-optimal solution via Gaussian randomization. We follow the rand C method and generate a set $\mathcal{R}$ of randomized solutions denoted by $\hat{\boldsymbol{\theta}}^{\text{rd}}$. We normalize them as $\boldsymbol{\eta}(\hat{\boldsymbol{\theta}}^{\text{rd}})=\hat{\boldsymbol{\theta}}^{\text{rd}}/\hat{\theta}^{\text{rd}}_{N+1}$ to ensure that the last element is one and then we define $\tilde{\eta}(\hat{\boldsymbol{\theta}}^{\text{rd}})_{n}=e^{j\angle\eta(\hat{\boldsymbol{\theta}}^{\text{rd}})_{n}},\leavevmode\nobreak\ n=1,\dots,N$, to meet the unit modulus constraint. We then define the set of feasible randomized solutions as =$\mathcal{E}=\Big{\{}\boldsymbol{\tilde{\eta}}\left(\hat{\boldsymbol{\theta}}^{\text{rd}}\right):\hat{\boldsymbol{\theta}}^{\text{rd}}\in\mathcal{R},\leavevmode\nobreak\ \text{SINR}_{k}\left(\boldsymbol{\tilde{\eta}}\left(\hat{\boldsymbol{\theta}}^{\text{rd}}\right),\mathbf{F}^{\text{opt}}\right)\geq\gamma_{k},\leavevmode\nobreak\ k=1,\dots,K\Big{\}}$, the RIS phase-shifts are recovered as
The steps of the AO algorithm are outlined in Algorithm 1.
@@
4:   Retrieve $\mathbf{F}^{(t+1)}$ trough Gaussian randomization
@@
6:   Retrieve $\boldsymbol{\theta}^{(t+1)}$ trough Gaussian randomization
@@ IV Malicious RIS sensing SNR degradation
As mentioned in the introduction, we assume that the RIS control circuit has been hacked. The hacker aims at degrading the sensing SNR so that the sensing target can go undetected. Furthermore, to make the hacker’s action hard to detect, the UEs are still guaranteed a minimum SINR. Motivated by the collaboration between the BS and RIS, we assume that the hacker has access to the optimal precoders $\mathbf{F}^{\text{opt}}$. Additionally, we consider the worst-case scenario in which perfect CSI is available to the hacker. Hence, the hacker models the cascaded channel as $\mathbf{h}_{k}=\mathbf{h}_{s,k}+\mathbf{H}_{t}^{\mathsf{H}}\mathbf{H}_{r,k}\boldsymbol{\theta}$. According to these assumptions, the sensing SNR per stream can be rewritten as
@@
$\quad+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{M}_{l}\boldsymbol{\theta}+\delta_{s}^{2}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{s,{\rm T}}^{\mathsf{H}}\mathbf{H}_{s,{\rm T}}\mathbf{f}_{l},$ (23)
@@
$\mathbf{M}_{l}$ $=\mathbf{G}^{\mathsf{H}}\mathbf{H}_{t}\left(\delta_{m}^{*}\mathbf{H}_{s,{\rm T}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}+\delta_{m}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{s,{\rm T}}^{\mathsf{H}}\right)\mathbf{H}_{t}^{\mathsf{H}}\mathbf{G}.$ (25)
@@
$\text{P}3:=\underset{\boldsymbol{\theta}}{\textrm{minimize}}$ $\leavevmode\nobreak\ \frac{1}{\sigma_{\rm T}^{2}}\sum_{l=1}^{L}\rho_{l}$ (26a)
subject to $\text{SINR}_{k}(\boldsymbol{\theta},\mathbf{F}^{\text{opt}})\geq\gamma_{k},\quad k=1,\dots,K,$ (26b)
@@
This problem is non-convex and therefore challenging to solve. However, we will demonstrate that each non-convex term of the objective function can be reformulated in a manageable manner, enabling us to solve P3333 using the convex-concave procedure (CCP) [13]. The first term of $\rho_{l}$ is the product between individually convex terms of $\boldsymbol{\theta}$. We decouple those terms by introducing the variables $\mathbf{s}=[s_{1}\dots s_{L}]^{\top}$ and $r$, redefining $\rho_{l}$ by removing the constant term as
$\rho_{l}$ $=\delta_{r}^{2}s_{l}^{2}r^{2}+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{M}_{l}\boldsymbol{\theta}$ (27)
and adding the constraints
$|\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\mathbf{G}\boldsymbol{\theta}|^{2}\leq s_{l}^{2},\quad l=1,\dots,L,$ (28)
$\|\mathbf{H}_{t}^{\mathsf{H}}\mathbf{G}\boldsymbol{\theta}\|^{2}\leq r^{2}.$ (29)
@@
where $\hat{\mathbf{M}}_{l}^{(+)},\hat{\mathbf{M}}_{l}^{(-)}$ are respectively the positive and negative semi-definite parts of $\hat{\mathbf{M}}_{l}$ and $\hat{\boldsymbol{\theta}}^{(t)}$ being a local point. The second and third term of (30) are concave and are then locally upper bounded by their first-order Taylor expansion. We define an approximated version of $\rho_{l}$ around the local point $(\mathbf{s}^{(t)},r^{(t)},\hat{\boldsymbol{\theta}}^{(t)})$ (by also removing the constant terms):
@@
6:until $\left\|\boldsymbol{\theta}^{(t)}-\boldsymbol{\theta}^{(t-1)}\right\|\leq\nu,\|\boldsymbol{\xi}\|\leq\nu,\|\mathbf{v}\|\leq\nu$
@@
Algorithm 2 Sensing SNR minimization algorithm
where $\boldsymbol{\xi}=[\xi_{1}\dots\xi_{2N}]^{\top}$ is a vector of auxiliary variables. The minimum SINR constraint (26b) is written as:
$\gamma_{k}\left(\hat{\boldsymbol{\theta}}^{\mathsf{H}}\mathbf{U}_{k}\hat{\boldsymbol{\theta}}+\sigma_{k}^{2}\right)-\hat{\boldsymbol{\theta}}^{\mathsf{H}}\mathbf{R}_{k,k}\hat{\boldsymbol{\theta}}\leq 0$ (36)
where
$\mathbf{U}_{k}$ $=\begin{bmatrix}\mathbf{H}_{r,k}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{F}_{-k}\mathbf{F}_{-k}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\mathbf{H}_{r,k}&\mathbf{H}_{r,k}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{F}_{-k}\mathbf{F}_{-k}^{\mathsf{H}}\mathbf{h}_{s,k}\\ \mathbf{h}_{s,k}^{\mathsf{H}}\mathbf{F}_{-k}\mathbf{F}_{-k}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\mathbf{H}_{r,k}&\mathbf{h}_{s,k}^{\mathsf{H}}\mathbf{F}_{-k}\mathbf{F}_{-k}^{\mathsf{H}}\mathbf{h}_{s,k}\end{bmatrix}.$ (37)
The notation $\mathbf{F}_{-k}\in\mathbb{C}^{M\times L-1}$ indicates a matrix containing all the precoding vectors except the one assigned to the $k$-th UE. (36) is then convexified as
where $\mathbf{v}=[v_{1}\dots v_{K}]^{\top}$ are slack variables. Lastly, (28) and (29) are locally approximated as
$\text{P}4:=\underset{\hat{\boldsymbol{\theta}},\mathbf{v},\boldsymbol{\xi},\mathbf{s},r}{\textrm{minimize}}$ $\leavevmode\nobreak\ \frac{1}{\sigma_{\rm T}^{2}}\sum_{l=1}^{L}\tilde{\rho}_{l}^{(t)}+\lambda^{(t)}(||\boldsymbol{\xi}||+||\mathbf{v}||)$ (41a)
subject to $\hat{\theta}_{N+1}=1,$ (41b)
[…]
$\boldsymbol{\xi},\mathbf{v},\mathbf{s},r\geq 0,$ (41d)

### v3

@@ III-B RIS phase-shift optimization
We now move on to the second part of the AO, where we obtain $\boldsymbol{\Theta}$ while $\mathbf{F}$ is fixed. We rewrite $\rho_{l}$, defined in (6), as
@@
$\quad+2\Re\left(\text{Tr}\left(\delta_{m}^{*}\boldsymbol{\Theta}\mathbf{T}^{\mathsf{H}}\boldsymbol{\Theta}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{h}_{\textrm{s,T}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\right)\right)$
$\quad+\delta_{s}^{2}\text{Tr}\left(\mathbf{h}_{\textrm{s,T}}^{\mathsf{H}}\mathbf{h}_{\textrm{s,T}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\right).$ (12)
By utilizing the identity $\text{Tr}(\mathbf{A}\mathbf{X}\mathbf{B}\mathbf{X})=(\text{vec}\left(\mathbf{X}^{\top}\right)^{\top})(\mathbf{B}^{\top}\otimes\mathbf{A})\text{vec}(\mathbf{X})$ and $\mathbf{T}=\mathbf{T}^{\mathsf{H}}$, the first term of $\rho_{l}$ can be rewritten as
$\text{vec}\left(\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}}\right)^{\mathsf{H}}\left(\left(\mathbf{H}_{t}^{*}\mathbf{H}_{t}^{\top}\right)\otimes\left(\delta_{r}^{2}\mathbf{H}_{\text{t}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\right)\right)\text{vec}\left(\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}}\right)$ (13)
Then, thanks to the nature of $\boldsymbol{\Theta}$ [10], we can write that $\text{vec}(\boldsymbol{\Theta}\mathbf{T}\boldsymbol{\Theta}^{\mathsf{H}})^{\mathsf{H}}=\text{diag}(\text{vec}(\mathbf{T}))^{\mathsf{H}}\tilde{\boldsymbol{\theta}}$, where $\tilde{\boldsymbol{\theta}}$ collects the diagonal entries of $\boldsymbol{\Theta}^{*}\otimes\boldsymbol{\Theta}$. We can finally rewrite the first term of $\rho_{l}$ as $\tilde{\boldsymbol{\theta}}^{\mathsf{H}}\mathbf{Q}_{l}\tilde{\boldsymbol{\theta}}$, where
$\mathbf{Q}_{l}$ $=\left(\text{diag}\left(\text{vec}\left(\mathbf{T}\right)\right)\right)^{\mathsf{H}}\left(\left(\mathbf{H}_{t}^{*}\mathbf{H}_{t}^{\top}\right)^{\top}\otimes\left(\delta_{r}^{2}\mathbf{H}_{\text{t}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\right)\right)$
@@
$\mathbf{p}_{l}$ $=\text{diag}(\text{vec}(\mathbf{T}))^{\mathsf{H}}\text{vec}\left(\delta_{m}^{*}\mathbf{H}_{t}\mathbf{h}_{\textrm{s,T}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\right).$ (16)
@@
$=\frac{1}{\sigma_{\rm T}^{2}}\sum_{l=1}^{L}2\Re\left(\boldsymbol{\theta}^{\mathsf{H}}\left(\boldsymbol{\Sigma}_{\mathbf{b}_{l}^{(t)}}+\boldsymbol{\Sigma}_{\mathbf{p}_{l}}\right)\boldsymbol{\theta}\right),$ (18)
where $\boldsymbol{\Sigma}_{\mathbf{b}_{l}^{(t)}}\in\mathbb{C}^{N\times N}$ is a matrix built by stacking the elements of $\mathbf{b}_{l}^{(t)}$. We define the variable $\hat{\boldsymbol{\Theta}}=\hat{\boldsymbol{\theta}}\hat{\boldsymbol{\theta}}^{\mathsf{H}}$, where $\hat{\boldsymbol{\theta}}=[\boldsymbol{\theta}^{\top},1]^{\top}$ and $\mathbf{h}_{\textrm{r},k}=\text{diag}(\mathbf{h}_{\textrm{r},k})$. Then, the second subproblem of the AO algorithm is
@@
$\text{Tr}\left(\mathbf{R}_{k,k}\hat{\boldsymbol{\Theta}}\right)\geq\gamma_{k}\left(\sum_{l\neq k}\text{Tr}\left(\mathbf{R}_{k,l}\hat{\boldsymbol{\Theta}}\right)+\sigma^{2}_{k}\right)$ (19c)
@@
4:   Retrieve $\mathbf{F}^{(t+1)}$ through Gaussian randomization
@@
6:   Retrieve $\boldsymbol{\theta}^{(t+1)}$ through Gaussian randomization
@@
where
$\mathbf{R}_{k,l}$ $=\begin{bmatrix}\mathbf{h}_{\textrm{r},k}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\mathbf{h}_{\textrm{r},k}&\mathbf{h}_{\textrm{r},k}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{h}_{\textrm{s},k}\\ \mathbf{h}_{\textrm{s},k}^{\mathsf{H}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{\text{t}}^{\mathsf{H}}\mathbf{h}_{\textrm{r},k}&|\mathbf{h}_{\textrm{s},k}^{\mathsf{H}}\mathbf{f}_{l}|^{2}\end{bmatrix},$ (20)
$\boldsymbol{\Xi}_{l}^{(t)}$ $=\begin{bmatrix}\boldsymbol{\Sigma}_{\mathbf{b}_{l}^{(t)}}+\boldsymbol{\Sigma}_{\mathbf{p}_{l}}&\mathbf{0}_{N\times 1}\\ \mathbf{0}_{1\times N}&0\end{bmatrix}.$ (21)
By removing the rank-one constraints, the relaxed problem becomes convex. Hence, it can be solved by CVX, and we can retrieve a good sub-optimal solution via Gaussian randomization. Following the rand C method, we generate a set $\mathcal{R}$ of randomized solutions denoted by $\hat{\boldsymbol{\theta}}^{\text{rd}}$. We normalize them as $\boldsymbol{\eta}(\hat{\boldsymbol{\theta}}^{\text{rd}})=\hat{\boldsymbol{\theta}}^{\text{rd}}/\hat{\theta}^{\text{rd}}_{N+1}$ to ensure that the last element is one and then we define $\tilde{\eta}(\hat{\boldsymbol{\theta}}^{\text{rd}})_{n}=e^{j\angle\eta(\hat{\boldsymbol{\theta}}^{\text{rd}})_{n}},\leavevmode\nobreak\ n=1,\dots,N$, to meet the unit modulus constraint. We then define the set of feasible randomized solutions as =$\mathcal{E}=\Big{\{}\boldsymbol{\tilde{\eta}}\left(\hat{\boldsymbol{\theta}}^{\text{rd}}\right):\hat{\boldsymbol{\theta}}^{\text{rd}}\in\mathcal{R},\leavevmode\nobreak\ \text{SINR}_{k}\left(\boldsymbol{\tilde{\eta}}\left(\hat{\boldsymbol{\theta}}^{\text{rd}}\right),\mathbf{F}^{\text{opt}}\right)\geq\gamma_{k},\leavevmode\nobreak\ k=1,\dots,K\Big{\}}$, the RIS phase-shifts are recovered as
The steps of the AO algorithm are outlined in Algorithm 1.
@@ IV Malicious RIS sensing SNR degradation
We assume that the RIS control circuit has been hacked. The hacker aims to degrade the sensing SNR so that the sensing target can go undetected. To make the hacker’s action hard to detect, the UEs are still guaranteed a minimum SINR. Motivated by the collaboration between the BS and RIS, we assume that the hacker has access to the optimal precoders $\mathbf{F}^{\text{opt}}$. Additionally, we consider the worst-case scenario in which perfect CSI is available to the hacker. Hence, the hacker models the cascaded channel as $\mathbf{h}_{k}=\mathbf{h}_{\textrm{s},k}+\mathbf{H}_{t}^{\mathsf{H}}\mathbf{h}_{\textrm{r},k}\boldsymbol{\theta}$. According to these assumptions, the $\rho_{l}$ can be rewritten as
@@
$\quad+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{M}_{l}\boldsymbol{\theta}+\delta_{s}^{2}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{h}_{\textrm{s,T}}^{\mathsf{H}}\mathbf{h}_{\textrm{s,T}}\mathbf{f}_{l},$ (23)
@@
$\mathbf{M}_{l}$ $=\mathbf{G}^{\mathsf{H}}\mathbf{H}_{t}\left(\delta_{m}^{*}\mathbf{h}_{\textrm{s,T}}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}+\delta_{m}\mathbf{f}_{l}\mathbf{f}_{l}^{\mathsf{H}}\mathbf{h}_{\textrm{s,T}}^{\mathsf{H}}\right)\mathbf{H}_{t}^{\mathsf{H}}\mathbf{G}.$ (25)
@@
$\text{P}3:=\underset{\boldsymbol{\theta}}{\textrm{minimize}}$ $\leavevmode\nobreak\ \frac{\sum_{l=1}^{L}\rho_{l}}{\sigma_{\rm T}^{2}}$ (26a)
subject to $\text{SINR}_{k}(\boldsymbol{\theta},\mathbf{F}^{\text{opt}})\geq\gamma_{k},\leavevmode\nobreak\ k=1,\dots,K,$ (26b)
@@
This problem is non-convex and, therefore, challenging to solve. However, we will demonstrate that each non-convex term of the objective function can be reformulated in a manageable manner, enabling us to solve P3333 using the convex-concave procedure (CCP) [13]. The first term of $\rho_{l}$ is the product between individually convex terms of $\boldsymbol{\theta}$. We decouple those terms by introducing the variables $\mathbf{s}=[s_{1}\dots s_{L}]^{\top}$ and $r$, redefining $\rho_{l}$ by removing the constant term as $\rho_{l}=\delta_{r}^{2}s_{l}^{2}r^{2}+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{M}_{l}\boldsymbol{\theta},$ and adding the constraints
$|\mathbf{f}_{l}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\mathbf{G}\boldsymbol{\theta}|^{2}\leq s_{l}^{2}\|\mathbf{H}_{t}^{\mathsf{H}}\mathbf{G}\boldsymbol{\theta}\|^{2}\leq r^{2},\quad l=1,\dots,L.$ (27)
@@
where $\hat{\mathbf{M}}_{l}^{(+)},\hat{\mathbf{M}}_{l}^{(-)}$ are respectively the positive and negative semi-definite parts of $\hat{\mathbf{M}}_{l}$ and $\hat{\boldsymbol{\theta}}^{(t)}$ being a local point. The second and third terms of (28) are concave and are then locally upper bounded by their first-order Taylor expansion. We define an approximated version of $\rho_{l}$ around the local point $(\mathbf{s}^{(t)},r^{(t)},\hat{\boldsymbol{\theta}}^{(t)})$ by also removing the constant terms:
@@
where $\boldsymbol{\xi}=[\xi_{1}\dots\xi_{2N}]^{\top}$ is a vector of auxiliary variables. The minimum SINR constraint (26b) is written as
$\gamma_{k}\left(\hat{\boldsymbol{\theta}}^{\mathsf{H}}\mathbf{U}_{k}\hat{\boldsymbol{\theta}}+\sigma_{k}^{2}\right)-\hat{\boldsymbol{\theta}}^{\mathsf{H}}\mathbf{R}_{k,k}\hat{\boldsymbol{\theta}}\leq 0,$ (34)
where
$\mathbf{U}_{k}$ $=\begin{bmatrix}\mathbf{h}_{\textrm{r},k}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{F}_{-k}\mathbf{F}_{-k}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\mathbf{h}_{\textrm{r},k}&\mathbf{h}_{\textrm{r},k}^{\mathsf{H}}\mathbf{H}_{t}\mathbf{F}_{-k}\mathbf{F}_{-k}^{\mathsf{H}}\mathbf{h}_{\textrm{s},k}\\ \mathbf{h}_{\textrm{s},k}^{\mathsf{H}}\mathbf{F}_{-k}\mathbf{F}_{-k}^{\mathsf{H}}\mathbf{H}_{t}^{\mathsf{H}}\mathbf{h}_{\textrm{r},k}&\mathbf{h}_{\textrm{s},k}^{\mathsf{H}}\mathbf{F}_{-k}\mathbf{F}_{-k}^{\mathsf{H}}\mathbf{h}_{\textrm{s},k}\end{bmatrix}.$ (35)
The notation $\mathbf{F}_{-k}\in\mathbb{C}^{M\times L-1}$ indicates a matrix containing all the precoding vectors except the one assigned to the $k$-th UE. The constraint in (34) is then convexified as
where $\mathbf{v}=[v_{1}\dots v_{K}]^{\top}$ are auxiliary variables. Lastly, (27) are locally approximated as
$\text{P}4:=\underset{\hat{\boldsymbol{\theta}},\mathbf{v},\boldsymbol{\xi},\mathbf{s},r}{\textrm{minimize}}$ $\leavevmode\nobreak\ \frac{\sum_{l=1}^{L}\tilde{\rho}_{l}^{(t)}}{\sigma_{\rm T}^{2}}+\lambda^{(t)}(||\boldsymbol{\xi}||+||\mathbf{v}||)$ (39a)
subject to $\hat{\theta}_{N+1}=1,$ (39b)
[…]
$\boldsymbol{\xi},\mathbf{v},\mathbf{s},r\geq 0,$ (39d)
@@
6:until $\left\|\hat{\boldsymbol{\theta}}^{(t)}-\hat{\boldsymbol{\theta}}^{(t-1)}\right\|\leq\nu,\|\boldsymbol{\xi}\|\leq\nu,\|\mathbf{v}\|\leq\nu$
@@
Algorithm 2 Sensing SNR degradation algorithm